OpenStudy (anonymous):

Find the coordinates for the midpoint of the segment with endpoints given. (3, 5) and (-2, 0)

OpenStudy (anonymous):

cordinates of midpoint be P(x,y) \[P(x,y)=(\frac{x_1+x_2}{2}, \frac{y_1+y_2}{2})=(\frac{3+(-2)}{2}, \frac{5+0}{2})\] \[=(\frac{3-2}{2}, \frac{5}{2})=(\frac{1}{2}, \frac{5}{2})=(\frac{1}{2}, \frac{5}{2}) -or-(0.5, 2.5)\]